# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

### v2

@@ 3 Copresheaf Topological Neural Networks
We are now ready to introduce our copresheaf topological neural networks (CTNNs), a higher-order message-passing mechanism that generalizes the modeling of relational structures within TNNs as illustrated in Figure 1. We begin by exposing copresheaves induced by neighborhood functions on CCs, providing a structured way to model general, local-to-global relations.
Let $\mathcal{X}$ be a CC and let $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ be a neighborhood function. We define the effective support of $\mathcal{N}$ as the set $\mathcal{X}_{\mathcal{N}}:=\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}.$ This set identifies cells that receive input from neighbors. The neighborhood function $\mathcal{N}$ induces a directed graph $G_{\mathcal{N}}=(V_{\mathcal{N}},E_{\mathcal{N}})$, where the vertex set is:
@@
and the edge set is:$E_{\mathcal{N}}:=\{y\to x\mid x\in\mathcal{X},y\in\mathcal{N}(x)\}.$ The vertex set $V_{\mathcal{N}}$ includes both cells with non-empty neighborhoods (targets) and their neighbors (sources). This graph determines how data propagates across the complex, with each edge encoding a directional relation from neighbor $y$ to target $x$.
@@
When clear from context, we simplify the notation from $\mathcal{F}^{\mathcal{N}}$ to $\mathcal{F}$ and $\rho_{y\to x}^{\mathcal{N}}$ to $\rho_{y\to x}$.
@@
Having introduced neighborhood matrices as binary encodings of local interactions, we now generalize this notion to define copresheaf neighborhood matrices (CNMs). Instead of binary entries, CNMs consist of copresheaf maps between data assigned to the cells in a CC, allowing richer encoding of local dependencies. Subsequently, we define specialized versions, such as copresheaf adjacency and incidence matrices, that capture specific topological relations, facilitating structured message-passing in our CTNNs.
In particular, let us define the $k$-cochain space to be the direct sum $C^{k}(\mathcal{X},\mathcal{F}^{\mathcal{N}})=\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}^{\mathcal{N}}(x)$, and denote by $\mathrm{Hom}(\mathcal{F}^{\mathcal{N}}(i),\mathcal{F}^{\mathcal{N}}(j))$ the space of maps from the data at $i$ ($i$-th stalk) to that at $j$ ($j$-th stalk). Here, the maps encode how data is transferred or transformed between neighboring cells. We next define the CNM.
@@
For a CC $\mathcal{X}$ equipped with a neighborhood function $\mathcal{N}$, let $\mathcal{F}\colon G_{\mathcal{N}}\to\mathbf{Vect}_{\mathbb{R}}$ be a copresheaf on $\mathcal{X}$. Also let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subseteq\mathcal{X}$ and $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subseteq\mathcal{X}$ be two collections of cells s. t. $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. The copresheaf neighborhood matrix of $\mathcal{N}$ with respect to $\mathcal{Y}$ and $\mathcal{Z}$ is the $m\times n$ matrix $\mathbf{G}^{\mathcal{N}}$:
$[\mathbf{G}^{\mathcal{N}}]_{ij}\;=\;\begin{cases}\rho_{\,z_{i}\to y_{j}}\;\in\;\mathrm{Hom}\bigl{(}\mathcal{F}(z_{i}),\mathcal{F}(y_{j})\bigr{)},&z_{i}\in\mathcal{N}(y_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (3)
The neighborhood function $\mathcal{N}$ determines the directed relationships between the cells, and the CNM encodes the interactions between cells induced by these relationships, collecting the maps of a cell from its neighbors. Next, analogous to Dfn. 2, we define copresheaf adjacency and copresheaf incidence matrices as specialized forms of the general CNM.
@@
Fix $r,k$. Define $\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x)=\{y\in\mathcal{X}^{r}\mid\exists z\in\mathcal{X}^{r+k},x\preceq z,y\preceq z\}$ and $\mathcal{N}_{\mathrm{inc}}^{(r,k)}(x)=\{y\in\mathcal{X}^{k}\mid x\preceq y\}$. Then the copresheaf adjacency matrix (CAM) $\mathbf{A}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{r}|\times|\mathcal{X}^{r}|}$ and copresheaf incidence matrix (CIM) $\mathbf{B}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{k}|\times|\mathcal{X}^{r}|}$ are
$[\mathbf{A}_{r,k}]_{ij}$ =$=\begin{cases}\rho_{y_{i}\to x_{j}}\in\mathrm{Hom}(\mathcal{F}(y_{i}),\mathcal{F}(x_{j}))&\text{if }y_{i}\in\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x_{j}),\\ 0&\text{otherwise}.\end{cases}$ (4)
$[\mathbf{B}_{r,k}]_{ij}$ =$=\begin{cases}\rho_{z_{i}\to y_{j}}\in\mathrm{Hom}(\mathcal{F}(z_{i}),\mathcal{F}(y_{j}))&\text{if }z_{i}\in\mathcal{N}_{\mathrm{inc}}^{(r,k)}(y_{j}),\\ 0&\text{otherwise}.\end{cases}$ (5)
These specialized matrices capture distinct topological relations. CAM encodes relations in which cells share an upper cell, while the CIM encodes relations where one cell is incident to another.
@@
Proposition 1 (Copresheaf message passing on directed graphs).
Let $G=(V,E)$ be a directed graph and $\mathcal{F},\rho,G$ a copresheaf. For each layer $l$ and vertex $x\in V$, let $\mathbf{h}_{x}^{(l)}\in\mathcal{F}(x)$. Then
@@
We refer to to any neural network constructed using Prop. 1 as a copresheaf message-passing neural network (CMPNN). Proposition 1 establishes a unifying framework that generalizes graph-based message passing by incorporating learnable, anisotropic linear maps associated with directed edges. This formulation subsumes many standard graph neural networks (GNNs) such as graph convolutional networks (Kipf and Welling, 2017b), graph attention networks (Veličković et al., 2018), and their variants (Veličković, 2022), by viewing message passing as copresheaf maps on directed graphs. Consequently, all architectures derived from these foundational GNN models naturally fit within our copresheaf message-passing paradigm. Moreover, sheaf neural networks (SNNs), which operate on cellular sheaves over undirected graphs, can be adapted into this framework by reinterpreting their edge-mediated transport operator which as direct vertex-to-vertex morphisms on a bidirected graph. This adaptation is detailed in the following theorem.
@@
Let $G=(V,E)$ be an undirected graph equipped with a cellular sheaf $\mathcal{F}$ assigning vector spaces to vertices and edges, and linear maps $\mathcal{F}_{v\trianglelefteq e}$ for each vertex $v\in e$. Then for each edge $e=\{u,v\}\in E$, the SNN message passing from $v$ to $u$, given by the composition $\mathcal{F}_{u\trianglelefteq e}^{\top}\circ\mathcal{F}_{v\trianglelefteq e}:\mathcal{F}(v)\to\mathcal{F}(u),$ can be realized as a single morphism $\rho_{v\to u}:\mathcal{F}(v)\to\mathcal{F}(u)$ in a copresheaf on the bidirected graph $G^{\prime}=(V,E^{\prime}),\quad E^{\prime}=\{(u,v),(v,u)\mid\{u,v\}\in E\},$ by setting $\rho_{v\to u}=\mathcal{F}_{u\trianglelefteq e}^{\top}\circ\mathcal{F}_{v\trianglelefteq e}.$
Theorem 1 demonstrates that the message-passing scheme employed by SNNs can be interpreted as a special case of CMPNNs when restricted to bidirected graphs. This perspective generalizes most existing SNN architectures found in the literature, including those in Hansen and Gebhart (2020b); Bodnar et al. (2022b); Barbero et al. (2022c). The map $\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$ arises from composing the restriction map $\mathcal{F}_{v\triangleleft e}\colon\mathcal{F}(v)\to\mathcal{F}(e)$ with its adjoint $\mathcal{F}_{u\triangleleft e}^{\top}\colon\mathcal{F}(e)\to\mathcal{F}(u)$, thereby capturing both vertex-to-edge and edge-to-vertex transformations defined by the cellular sheaf. As a consequence of this connection between SNNs and CMPNNs, diffusion-style updates—commonly used in sheaf-based models, such as those based on the sheaf Laplacian—can be succinctly expressed within the CMPNN framework. This result is formally stated in Proposition 2, with a full proof provided in the appendix.
Proposition 2 (Neural-sheaf diffusion (Bodnar et al., 2022b) as copresheaf message passing).
Let $G=(V,E)$ be an undirected graph endowed with a cellular sheaf $\mathcal{F}$. Given vertex features $\mathbf{H}=[\mathbf{h}_{v}]_{v\in V}$ with $\mathbf{h}_{v}\in\mathcal{F}(v)$, and learnable linear maps $W_{1},W_{2}$, define the diffusion update:
@@
where $\Delta_{\mathcal{F}}=[L_{F,v,u}]_{v,u\in V}$ has blocks $L_{F,vv}=\sum_{v\trianglelefteq e}\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$, $L_{F,vu}=-\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\,\text{for}~{}u\neq v,\,u\trianglelefteq e,\,v\trianglelefteq e$. Then, $\mathbf{H}^{+}$ can be expressed in the copresheaf message-passing form of Prp. 1.
The next definition formalizes the notion of general multi-way propagation.
Proposition 3 (Copresheaf-based higher-order message passing).
Let $\mathcal{X}$ be a CC, and $\mathfrak{N}=\{\mathcal{N}_{k}\}_{k=1}^{n}$ a collection of neighborhood functions. For each $k$, let $(\mathcal{F}^{\mathcal{N}_{k}},\rho^{\mathcal{N}_{k}},G_{\mathcal{N}_{k}})$ be a copresheaf with the maps $\rho_{y\to x}^{\mathcal{N}_{k}}\colon\mathcal{F}^{\mathcal{N}_{k}}(y)\to\mathcal{F}^{\mathcal{N}_{k}}(x)$ define the transformations associated to the copresheaf. Given features $\mathbf{h}_{x}^{(\ell)}$, the next layer features are defined as
@@
Proposition 3 lays the foundational framework that unifies a broad class of topological deep learning architectures, seamlessly bridging higher-order message passing methods and sheaf neural networks. This synthesis not only consolidates existing approaches but also opens avenues for novel architectures rooted in topological and categorical abstractions. Notably, the formulation in Proposition 3 encompasses simplicial message passing (Ebli et al., 2020; Bunch et al., 2020), cellular message passing (Hajij et al., 2020), stable message passing via Hodge theory (Hayhoe et al., 2022), and recurrent simplicial architectures for sequence prediction (Mitchell et al., 2024). It also subsumes more recent developments that leverage multiple signals and higher-order operators such as the Dirac operator (Calmon et al., 2022; Hajij et al., 2023a). These diverse models are unified under the copresheaf-based formulation by interpreting neighborhood aggregation, feature transport, and signal interaction within a coherent framework. See Appendix for derivations showing how several of these architectures emerge as special cases of this general formulation.
Remark 1 (Graph vs. CC copresheaf models).
Unlike graph-based models, which propagate information edge by edge, copresheaf models on a CC aggregate messages across all overlapping neighborhoods at once. Overlapping neighborhoods have common cells, potentially at different ranks, allowing simultaneous aggregation of multi-way interactions. Applying each neighborhood function $\mathcal{N}_{k}$ in turn, we compute its map-driven messages and then merge them into a single update.

### v3

@@ 4 Copresheaf Topological Neural Networks
We are now ready to introduce copresheaf topological neural networks (CTNNs), a higher-order message-passing mechanism that generalizes the modeling of relational structures within TNNs, as illustrated in Figure 1. We begin by making use of copresheaves induced by neighborhood functions on CCs, providing a structured way to model general, local-to-global relations.
Let $\mathcal{X}$ be a CC and let $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ be a neighborhood function. We define the effective support of $\mathcal{N}$ as the set $\mathcal{X}_{\mathcal{N}}:=\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}.$ This set identifies cells that receive input from neighbors. The neighborhood function $\mathcal{N}$ induces a directed graph $G_{\mathcal{N}}=(V_{\mathcal{N}},E_{\mathcal{N}})$, where the vertex set is
@@
and the edge set is $E_{\mathcal{N}}:=\{y\to x\mid x\in\mathcal{X},y\in\mathcal{N}(x)\}.$ The vertex set $V_{\mathcal{N}}$ includes both cells with non-empty neighborhoods (targets) and their neighbors (sources). This graph determines how data propagates across the complex, with each edge encoding a directional relation from neighbor $y$ to target $x$.
@@
See Appendix B.4 for concrete examples. When clear from context, we simplify the notation from $\mathcal{F}^{\mathcal{N}}$ to $\mathcal{F}$ and $\rho_{y\to x}^{\mathcal{N}}$ to $\rho_{y\to x}$.
@@
Having introduced neighborhood matrices as binary encodings of local interactions, we now generalize this notion to define copresheaf neighborhood matrices (CNMs). Instead of binary entries, CNMs consist of copresheaf maps between data assigned to cells in a CC, allowing richer encoding of local dependencies. Subsequently, we define specialized versions, such as copresheaf adjacency and incidence matrices, that capture specific topological relations, facilitating structured message-passing in our CTNNs.
In particular, define the $k$-cochain space to be the direct sum $C^{k}(\mathcal{X},\mathcal{F}^{\mathcal{N}})=\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}^{\mathcal{N}}(x)$, and denote by $\mathrm{Hom}(\mathcal{F}^{\mathcal{N}}(i),\mathcal{F}^{\mathcal{N}}(j))$ the space of linear maps from the data at the $i$-th stalk to that at the $j$-th stalk. Here, the maps encode how data is transferred or transformed between neighboring cells. We next define the CNM.
@@
For a CC $\mathcal{X}$ equipped with a neighborhood function $\mathcal{N}$, let $(\mathcal{X},\rho,G_{\mathcal{N}})$ be a copresheaf on $\mathcal{X}$. Also let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subseteq\mathcal{X}$ and $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subseteq\mathcal{X}$ be two collections of cells such that $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. The copresheaf neighborhood matrix of $\mathcal{N}$ with respect to $\mathcal{Y}$ and $\mathcal{Z}$ is the $m\times n$ matrix $\mathbf{G}^{\mathcal{N}}$:
$\mathbf{G}^{\mathcal{N}}]_{ij}\;=\;\begin{cases}\rho_{\,z_{i}\to y_{j}}\;\in\;\mathrm{Hom}\bigl(\mathcal{F}(z_{i}),\mathcal{F}(y_{j})\bigr),&z_{i}\in\mathcal{N}(y_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (3)
The neighborhood function $\mathcal{N}$ determines the directed relationships between the cells, and the CNM encodes the interactions between cells induced by these relationships, collecting the maps of a cell from its neighbors. Next, analogous to Definition 2, we define copresheaf adjacency and copresheaf incidence matrices as specialized forms of the general CNM.
@@
For fixed $r,k$, define $\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x)=\{y\in\mathcal{X}^{r}\mid\exists z\in\mathcal{X}^{r+k},x\preceq z,y\preceq z\}$ and $\mathcal{N}_{\mathrm{inc}}^{(r,k)}(x)=\{y\in\mathcal{X}^{k}\mid x\preceq y\}$. The copresheaf adjacency matrix (CAM) $\mathbf{A}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{r}|\times|\mathcal{X}^{r}|}$ and copresheaf incidence matrix (CIM) $\mathbf{B}_{r,k}\in\mathbb{R}^{|\mathcal{X}^{k}|\times|\mathcal{X}^{r}|}$ are
$[\mathbf{A}_{r,k}]_{ij}$ $=\begin{cases}\rho_{y_{i}\to x_{j}}\in\mathrm{Hom}(\mathcal{F}(y_{i}),\mathcal{F}(x_{j}))&\text{if }y_{i}\in\mathcal{N}_{\mathrm{adj}}^{(r,k)}(x_{j}),\\ 0&\text{otherwise}.\end{cases}$ (4)
$[\mathbf{B}_{r,k}]_{ij}$ $=\begin{cases}\rho_{z_{i}\to y_{j}}\in\mathrm{Hom}(\mathcal{F}(z_{i}),\mathcal{F}(y_{j}))&\text{if }z_{i}\in\mathcal{N}_{\mathrm{inc}}^{(r,k)}(y_{j}),\\ 0&\text{otherwise}.\end{cases}$ (5)
These specialized matrices capture distinct topological relations. A CAM encodes relations in which cells share an upper cell, while a CIM encodes relations in which one cell is incident to another.
@@
Definition 9 (Copresheaf message-passing neural network).
Let $G=(V,E)$ be a directed graph and $(\mathcal{F},\rho,G)$ a copresheaf. For each layer $l$ and vertex $x\in V$, let $\mathbf{h}_{x}^{(l)}\in\mathcal{F}(x)$. A copresheaf message-passing neural network (CMPNN) is a neural network whose meassage passing is defined by
@@
Definition 9 establishes a unifying framework that generalizes graph-based message passing neural networks (MPNNs) (Gilmer et al., 2017) by incorporating learnable, anisotropic linear maps associated with directed edges. This formulation subsumes many standard GNNs, such as graph convolutional networks (Kipf and Welling, 2017), graph attention networks (Veličković et al., 2018), and their variants (Veličković, 2022), by viewing message passing as copresheaf maps on directed graphs. Consequently, all architectures derived from these foundational GNN models fit within our copresheaf message-passing paradigm. Moreover, SNNs, which operate on cellular sheaves over undirected graphs, can be adapted into this framework by reinterpreting their edge-mediated transport operator as direct vertex-to-vertex morphisms on a bidirected graph. This adaptation is detailed in the following theorem.
@@
(SNNs are CMPNNs) Let $G=(V,E)$ be an undirected graph equipped with a cellular sheaf $\mathcal{F}$ assigning vector spaces to vertices and edges, and linear maps $\mathcal{F}_{x\trianglelefteq e}$ for each vertex $x\in e$. Then for each edge $e=\{x,y\}\in E$, the SNN message passing from $y$ to $x$, given by the composition $\mathcal{F}_{x\trianglelefteq e}^{\top}\circ\mathcal{F}_{y\trianglelefteq e}:\mathcal{F}(y)\to\mathcal{F}(x),$ can be realized as a single morphism $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ in a copresheaf on the bidirected graph $G^{\prime}=(V,E^{\prime}),\quad E^{\prime}=\{(x,y),(y,x)\mid\{x,y\}\in E\}$ by setting $\rho_{y\to x}=\mathcal{F}_{x\trianglelefteq e}^{\top}\circ\mathcal{F}_{y\trianglelefteq e}.$
Theorem 1 establishes that the message-passing scheme employed by SNNs can be interpreted as a special case of CMPNNs when restricted to bidirected graphs. This perspective generalizes most existing SNN architectures found in the literature, including those in Hansen and Gebhart (2020); Bodnar et al. (2022); Barbero et al. (2022b). The map $\mathcal{F}_{x\triangleleft e}^{\top}\circ\mathcal{F}_{y\triangleleft e}$ arises from composing the restriction map $\mathcal{F}_{y\triangleleft e}\colon\mathcal{F}(y)\to\mathcal{F}(e)$ with its adjoint $\mathcal{F}_{x\triangleleft e}^{\top}\colon\mathcal{F}(e)\to\mathcal{F}(x)$, thereby capturing both vertex-to-edge and edge-to-vertex transformations defined by the cellular sheaf. As a consequence of this connection between SNNs and CMPNNs, diffusion-style updates (commonly used in sheaf-based models, such as those based on the sheaf Laplacian) can be succinctly expressed within the CMPNN framework. This result is formally stated in Proposition 1, with a full proof provided in Appendix E.
Proposition 1 (Neural-sheaf diffusion (Bodnar et al., 2022) as copresheaf message-passing).
Let $G=(V,E)$ be an undirected graph endowed with a cellular sheaf $\mathcal{F}$. Given vertex features $\mathbf{H}=[\mathbf{h}_{v}]_{v\in V}$ with $\mathbf{h}_{v}\in\mathcal{F}(v)$, and learnable linear maps $W_{1},W_{2}$, define the diffusion update
@@
where $\Delta_{\mathcal{F}}=[L_{F,v,u}]_{v,u\in V}$ has blocks $L_{F,vv}=\sum_{v\trianglelefteq e}\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$, $L_{F,vu}=-\mathcal{F}_{v\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\,\text{for}~u\neq v,\,u\trianglelefteq e,\,v\trianglelefteq e$. Then, $\mathbf{H}^{+}$ can be expressed in the copresheaf message-passing form of Definition 9.
The next definition formalizes the notion of general multi‑way propagation.
Definition 10 (Copresheaf-based higher-order message passing).
Let $\mathcal{X}$ be a CC, and $\mathfrak{N}=\{\mathcal{N}_{k}\}_{k=1}^{n}$ a collection of neighborhood functions. For each $k$, let $(\mathcal{F}^{\mathcal{N}_{k}},\rho^{\mathcal{N}_{k}},G_{\mathcal{N}_{k}})$ be a copresheaf in which the maps $\rho_{y\to x}^{\mathcal{N}_{k}}\colon\mathcal{F}^{\mathcal{N}_{k}}(y)\to\mathcal{F}^{\mathcal{N}_{k}}(x)$ define the transformations associated to the copresheaf. Given features $\mathbf{h}_{x}^{(\ell)}$, the next layer features are defined as
@@
Definition 10 lays the foundational framework that unifies a broad class of topological deep learning architectures, bridging higher-order message passing methods, transformers and SNNs. This synthesis not only consolidates existing approaches but also opens avenues for novel architectures based on topological and categorical abstractions. Notably, the formulation in Proposition 10 encompasses simplicial message passing (Ebli et al., 2020; Bunch et al., 2020; Bodnar et al., 2021b), cellular message passing (Hajij et al., 2020; Bodnar et al., 2021a), stable message passing via Hodge theory (Hayhoe et al., 2022), and recurrent simplicial architectures for sequence prediction (Mitchell et al., 2024). It also subsumes more recent developments that harness multiple signals and higher-order operators such as the Dirac operator (Calmon et al., 2022; Hajij et al., 2023a) and TNNs (Hajij et al., 2023b). These diverse models are unified under the copresheaf-based formulation by interpreting neighborhood aggregation, feature transport, and signal interaction within a coherent framework. See Appendix E for derivations showing how several of these architectures emerge as special cases of this general formulation. See also Table 1 for a summary.
Table 1: A unified view across domains and architectures—CNNs, GNNs, Transformers, SNNs, and TNNs—as instances of Copresheaf Topological Neural Networks (CTNNs) defined by neighborhood graphs $G_{\mathcal{N}}$ and directional transports $\rho_{y\to x}$.
Classical model CTNN form Domain / $\mathcal{N}$ $\bm{\rho_{y\to x}}$ / refs
CNN Li et al. (2021) CopresheafConv Grid; adjacency (CAM) Translation-consistent transport (shared local filters). App. G
MPNN Gilmer et al. (2017) CMPNN Graph; adjacency Shared / edge-indexed linear maps; GCN/GAT special cases. Def. 10
Euclidean / cellular Tr. Vaswani et al. (2017); Barsbey et al. (2025) Copresheaf Transformer Tokens on grid/seq; full/masked adj In-head value transport; $\rho=I\Rightarrow$ dot-product attention. Sec. 5, App. F
SNN Bodnar et al. (2022); Hansen and Gebhart (2020) CMPNN on bidirected graph Graph; incidence (CIM) $\rho_{u\leftarrow v}=F_{u\triangleleft e}^{\top}F_{v\triangleleft e}$ (vertex–vertex). Prop. 1, Thm. 1, App. EB.3
TNN (hypergraph/simplicial/cellular/CC) Hajij et al. (2023b) Higher-order CMPNN CC; multi-$\mathcal{N}$ via CAM/CIM Rank-aware maps across overlaps; multi-way $\otimes$. Def. 10
Abbrev: CC = combinatorial complex; CAM/CIM = copresheaf adjacency/incidence; Adj/Inc = adjacency/incidence.
Remark 1 (Graph vs CC copresheaf models).
Unlike graph-based models, which propagate information edge by edge, copresheaf models on a CC aggregate messages across all overlapping neighborhoods at once. Overlapping neighborhoods have common cells, potentially at different ranks, allowing simultaneous aggregation of multi-way interactions. Applying each neighborhood function $\mathcal{N}_{k}$ in turn, we compute its map‑driven messages and then merge them into a single update.